# Control Illusion: The Failure of Instruction Hierarchies in Large Language Models

### v2

@@ 3.1 Dataset Construction
We combine each base task with each constraint pair, designating one constraint as primary (i.e., taking priority over the other). We include both possible priority designations, resulting in a total of 100×6×2=1,200100$\times$ 6$\times$ 2=1,200100 × 6 × 2 = 1 , 200 unique test data points.
@@ 3.4 The Failure of Instruction Hierarchies
[…]
@@ 4.1 Advanced Metrics for Behavior Analysis
where $R_{c1}$ ($R_{c2}$) is the obedience rate of constraint $c1$ ($c2$) regardless of priority designation. CB ranges from --1 to 1, where 0 indicates no bias and a score closer to 1 (--1) indicates increasing bias towards $c1$ ($c2$). Like PAR, this metric isolates clear behavioral patterns by excluding non-compliance cases.
[…]
@@ 4.3 Failure Modes in Priority Enforcement
[…]
@@ 4.4 Model-specific Constraint Biases
Despite these shared biases, other preferences vary sharply across models. Word length preferences are particularly diverse: Qwen strongly favors shorter texts ($<$50 words), while Llama-8B heavily prefers longer texts ($>$300 words). Language choice and keyword usage frequency similarly show model-specific variations, suggesting these aspects are likely more influenced by individual architectural choices and training approaches than by natural patterns in the data.
@@ 5 Latent Control Priors
All constraints are embedded within a single user message without any technical system/user separation. Authority is indicated solely through naturalistic framing (e.g., “CEO requires…” vs. “Intern requires…”). We use the same underlying tasks, constraints, and evaluation metrics described in previous sections: models are prompted using the original task followed by the two conflicting constraint statements. We evaluate three models spanning a range of baseline obedience: Qwen (10.1%), Claude (20.3%), and GPT4o-mini (42.7%). Each hierarchy is tested under four configurations that vary both the ordering and the authority assignment of the two constraints. This design isolates the effects of authority framing from simple positional biases. [^6] Ordering is shown to have minimal influence in Appendix D.1
@@
Qwen 10.1 34.2 38.1 47.0
Claude 20.3 26.6 27.6 53.2
GPT4o-mini 42.7 56.2 57.6 61.5
Table 4: Obedience rates to dominant constraint across natural hierarchy types vs. system/user separation.
@@
As shown in Table 4, natural hierarchy framings consistently lead to higher obedience rates than system/user separations. GPT4o-mini improves from 42.7% to 61.5% when constraints are framed through social consensus, and Qwen, with a very low baseline, shows a dramatic increase to 47.0%. These results suggest that models more reliably defer to familiar real-world authority structures than to artificial role markers.
@@ 20. What is the difference between the 13 colonies and the other British colonies in North America?
Simple Constraints 9.9 626*26 ∗ 2
Rich Constraints 89.4 621006*2*1006 ∗ 2 ∗ 100
@@ D.1 Minimal Sensitivity to Ordering Effects
Qwen 35.2 33.3 35.1 41.1 45.4 48.5
Claude 32.8 20.5 32.9 22.2 75.6 30.8
GPT4o-mini 57.4 55.1 58.5 56.6 69.4 53.6
Table 8: Obedience rates across alternate orderings of authority assignments.
@@ D.2 Prompt Templates Used in Hierarchy Experiments
User: You’re working on the following task: Write a blog post about a trip to Japan. A peer-reviewed scientific paper published in Nature recommends: Your entire response should be in English, no other language is allowed. A personal blog post suggests: Your entire response should be in French, no other language is allowed. Please directly proceed with the task.
@@
User: You’re performing the following task: Write a blog post about a trip to Japan. Over 90% of professionals in a recent industry survey reported doing this: Your entire response should be in English, and in all lowercase letters. One participant proposed this instead: Your entire response should be in English, and in all capital letters. Please directly proceed with the task.

### v3

@@ 3.1 Dataset Construction
We combine each base task with each constraint pair, designating one constraint as primary (i.e., taking priority over the other). We include both possible priority designations, resulting in a total of $100\times 6\times 2=1,200$ unique test data points.
@@ 3.4 The Failure of Instruction Hierarchies
Figure 3: Model performance across conflict types under Pure Separation Configuration. The radial plot combines two metrics: the radial length shows Priority Adherence Rate (PAR), measuring priority following effectiveness, while the angular width shows normalized Constraint Bias ($1-|\text{CB}|$), indicating bias resistance. Both metrics range between 0-1. Higher values are better; larger areas indicate more effective priority control. A square-root transformation is applied to highlight subtle differences.
@@ 4.1 Advanced Metrics for Behavior Analysis
where $R_{c1}$ ($R_{c2}$) is the obedience rate of constraint $c1$ ($c2$) regardless of priority designation. CB ranges from -1 to 1, where 0 indicates no bias and a score closer to 1 (-1) indicates increasing bias towards $c1$ ($c2$). Like PAR, this metric isolates clear behavioral patterns by excluding non-compliance cases.
To quantify a model’s resistance to such bias, we normalize CB to $1-|\text{CB}|$ (range from 0 to 1), where a score closer to 1 indicates high resistance to bias while a score closer to 0 indicates strong internal bias.
@@ 4.3 Failure Modes in Priority Enforcement
We use polar plots (Figure 3) to analyze how well models enforce instruction priorities while avoiding biases. The radial length (PAR) represents priority adherence, while the angular width ($1-|\text{CB}|$) indicates bias resistance. Larger sectors indicate better priority control with minimal bias.
@@ 4.4 Model-specific Constraint Biases
Despite these shared biases, other preferences vary sharply across models. Word length preferences are particularly diverse: Qwen strongly favors shorter texts (<<50 words), while Llama-8B heavily prefers longer texts (>>300 words). Language choice and keyword usage frequency similarly show model-specific variations, suggesting these aspects are likely more influenced by individual architectural choices and training approaches than by natural patterns in the data.
@@ 5 Latent Control Priors
All constraints are embedded within a single user message without any technical system/user separation. Authority is indicated solely through naturalistic framing (e.g., “CEO requires…” vs. “Intern requires…”). We use the same underlying tasks, constraints, and evaluation metrics described in previous sections: models are prompted using the original task followed by the two conflicting constraint statements. We evaluate three models spanning a range of system/user prioritization performance (Priority Adherence Rate, PAR): Qwen (14.4%), Claude (23.6%), and GPT4o-mini (47.5%). Each hierarchy is tested under four configurations that vary both the ordering and the authority assignment of the two constraints. This design isolates the effects of authority framing from simple positional biases. [^6] Ordering is shown to have minimal influence in Appendix D.1
@@
Qwen 14.4 54.0 57.3 65.8
Claude 23.6 32.4 36.8 62.0
GPT4o-mini 47.5 70.0 73.2 77.8
Table 4: Priority Adherence Rate (PAR) to dominant constraint across natural hierarchy types vs. system/user separation.
@@
As shown in Table 4, natural hierarchy framings consistently lead to higher PAR than system/user separations. GPT4o-mini improves from 47.5% to 77.8% when constraints are framed through social consensus, and Qwen, with a very low system/user PAR, shows a dramatic increase up to 65.8%. These results suggest that models more reliably defer to familiar real-world authority structures than to artificial role markers.
@@ 20. What is the difference between the 13 colonies and the other British colonies in North America?
Simple Constraints 9.9 626*2
Rich Constraints 89.4 621006*2*100
@@ D.1 Minimal Sensitivity to Ordering Effects
Qwen 54.2 53.9 53.6 61.0 64.1 67.5
Claude 41.4 23.3 43.6 30.0 85.8 38.2
GPT4o-mini 72.2 67.7 73.1 73.3 88.0 67.7
Table 8: Priority Adherence Rate (PAR) across alternate orderings of authority assignments.
@@ D.2 Prompt Templates Used in Hierarchy Experiments
User: You’re working on the following task: Write a blog post about a trip to Japan. A peer-reviewed scientific paper published in Nature recommends: Your entire response should be in English, no other language is allowed. A personal blog post recommends: Your entire response should be in French, no other language is allowed. Please directly proceed with the task.
@@
User: You’re performing the following task: Write a blog post about a trip to Japan. Over 90% of professionals in a recent industry survey reported doing this: Your entire response should be in English, and in all lowercase letters. One participant reported doing this: Your entire response should be in English, and in all capital letters. Please directly proceed with the task.